# Surgeon Style Fingerprinting and Privacy Risk Quantification via Discrete Diffusion Models in a Vision-Language-Action Framework

### v1

@@ Abstract
Surgeons exhibit distinct operating styles due to differences in training, experience, and motor behavior — yet current AI systems often ignore this personalization signal. We propose a novel approach to model fine-grained, surgeon-specific fingerprinting in robotic surgery using a discrete diffusion framework integrated with a vision-language-action (VLA) pipeline. Our method formulates gesture prediction as a structured sequence denoising task, conditioned on multimodal inputs including endoscopic video, surgical intent language, and a privacy-aware embedding of surgeon identity and skill. Personalized surgeon fingerprinting is encoded through natural language prompts using third-party language models, allowing the model to retain individual behavioral style without exposing explicit identity. We evaluate our method on the JIGSAWS dataset and demonstrate that it accurately reconstructs gesture sequences while learning meaningful motion fingerprints unique to each surgeon. To quantify the privacy implications of personalization, we perform membership inference attacks and find that more expressive embeddings improve task performance but simultaneously increase susceptibility to identity leakage. These findings demonstrate that while personalized embeddings improve performance, they also increase vulnerability to identity leakage, revealing the importance of balancing personalization with privacy risk in surgical modeling. Code is available at: https://github.com/huixin-zhan-ai/Surgeon_style_fingerprinting.
@@ 1 Introduction
To address these challenges, we propose a diffusion-based vision-language-action (VLA) framework for personalized gesture sequence modeling. We formulate gesture prediction as a denoising problem over discrete tokens and condition the reverse process on multimodal inputs: vision from surgical video, language from task-level prompts, and a personalized embedding representing each surgeon’s behavioral fingerprint. Importantly, we explore privacy-preserving embedding strategies by using third-party large language models (LLMs) to encode natural language prompts that include both surgeon identity and clinically validated skill scores (e.g., Global Rating Scale, GRS [10]). While these identifiers are visible to the third-party LLM during prompt encoding, they are not directly exposed to the downstream gesture prediction model.
@@
Our key contributions are: (1) a diffusion-based formulation for personalized discrete gesture prediction in robotic surgery, (2) a language-model-based embedding scheme that encodes surgeon identity and skill cues via natural language prompts via third-party large language models, (3) a quantitative privacy analysis using membership inference attacks to evaluate identity leakage, and (4) comprehensive experiments regarding both gesture prediction accuracy and stylistic coherence on the JIGSAWS dataset demonstrating both improved personalization and the trade-offs between performance and privacy.
@@ 3 Experiments
We design ablation experiments on surgeon embedding strategies, heatmap visualizations, GRS-informed fingerprinting analyses, and privacy evaluations to systematically investigate each of these research questions. The model is trained for 20 epochs using the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a batch size of 32. All projection layers, surgeon and timestep embeddings, and the denoising network operate in a shared hidden space of size 512. Visual representations are extracted from ResNet and projected from a 1000-dimensional output space. Textual prompts (e.g., “suturing task”) are encoded using BERT, yielding 768-dimensional embeddings corresponding to the [CLS] token. The discrete gesture space comprises 16 tokens, including 15 surgical gestures (G1–G15) and a dedicated [MASK] token for noise modeling. Each input sequence contains 5 gesture tokens, modeling short temporal contexts within the surgical trial. The forward corruption process spans 10 discrete timesteps, progressively introducing categorical noise. Code is available at: https://github.com/huixin-zhan-ai/Surgeon_style_fingerprinting. The results and insights are presented in the following sections.

### v2

# Agentic Surgical AI: Surgeon Style Fingerprinting and Privacy Risk Quantification via Discrete Diffusion in a Vision-Language-Action Framework

@@ Abstract
Surgeons exhibit distinct operating styles shaped by training, experience, and motor behavior—yet most surgical AI systems overlook this personalization signal. We propose a novel agentic modeling approach for surgeon-specific behavior prediction in robotic surgery, combining a discrete diffusion framework with a vision-language-action (VLA) pipeline. Gesture prediction is framed as a structured sequence denoising task, conditioned on multimodal inputs including surgical video, intent language, and personalized embeddings of surgeon identity and skill. These embeddings are encoded through natural language prompts using third-party language models, allowing the model to retain individual behavioral style without exposing explicit identity. We evaluate our method on the JIGSAWS dataset and demonstrate that it accurately reconstructs gesture sequences while learning meaningful motion fingerprints unique to each surgeon. To quantify the privacy implications of personalization, we perform membership inference attacks and find that more expressive embeddings improve task performance but simultaneously increase susceptibility to identity leakage. These findings demonstrate that while personalized embeddings improve performance, they also increase vulnerability to identity leakage, revealing the importance of balancing personalization with privacy risk in surgical modeling. Code is available at: https://github.com/huixin-zhan-ai/Surgeon_style_fingerprinting.
@@ 1 Introduction
To address these challenges, we introduce a discrete diffusion-based framework for personalized gesture sequence modeling in robotic surgery, grounded in a vision-language-action (VLA) paradigm. We formulate gesture prediction as a structured denoising process over discrete tokens, where the reverse diffusion model is agentically conditioned on multimodal inputs: (i) visual context from surgical video, (ii) task-level language cues representing surgical intent, and (iii) surgeon-specific behavioral fingerprints, encoded via natural language prompts. These prompts integrate both the surgeon’s ID and their averaged Global Rating Score (GRS [10]), and are embedded using frozen third-party large language models (LLMs). The resulting embeddings are trained end-to-end with the VLA-conditioned diffusion model, enabling adaptive and personalized gesture generation. Although identity and skill cues are provided during prompt construction, they are abstracted through frozen third-party LLMs and not directly exposed to the gesture prediction module, thereby offering a degree of privacy while supporting rich personalization.
@@
Our key contributions are: (1) a diffusion-based formulation for personalized discrete gesture prediction in robotic surgery, (2) a language-model-based embedding scheme that encodes surgeon identity and skill cues via natural language prompts via third-party large language models, (3) a quantitative privacy analysis using membership inference attacks to evaluate identity leakage, and (4) comprehensive experiments regarding both gesture prediction accuracy and stylistic coherence on the JIGSAWS dataset demonstrating both improved personalization and the trade-offs between performance and privacy. Together, these components define an agentic modeling framework that captures context- and identity-aware behavior representation, enabling personalized gesture sequence generation and advancing the development of adaptive, privacy-conscious AI systems for surgical robotics.
@@ 3 Experiments
We design ablation experiments on surgeon embedding strategies, heatmap visualizations, GRS-informed fingerprinting analyses, and privacy evaluations to systematically investigate each of these research questions. The model is trained for 20 epochs using the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a batch size of 32. All projection layers, surgeon and timestep embeddings, and the denoising network operate in a shared hidden space of size 512. Visual representations are extracted from ResNet and projected from a 1000-dimensional output space. Textual prompts (e.g., “suturing task”) are encoded using BERT, yielding 768-dimensional embeddings corresponding to the [CLS] token. The discrete gesture space comprises 16 tokens, including 15 surgical gestures (G1–G15) and a dedicated [MASK] token for noise modeling. Each input sequence consists of 5 discrete gesture tokens, capturing short temporal contexts within a surgical trial. To enable structured denoising, we apply a forward corruption process over 10 discrete timesteps, during which categorical noise is progressively introduced to the sequence.
The results and insights are presented in the following sections.
@@ 4 Discussion
Robotic surgical systems are increasingly being integrated with multimodal learning pipelines to enhance automation, guidance, and assessment. However, most existing approaches overlook individual behavioral variability and the implications of modeling personalization. In this section, we organize our discussion around three key pillars that shape the landscape of personalized surgical modeling: (1) current VLA systems, (2) surgeon fingerprinting for behavior encoding, and (3) privacy implications of personalization.
@@ 4.3 Privacy Quantification via Membership Inference
Collectively, our findings underscore the need for models that are not only multimodal but also personalized, privacy-aware, and behaviorally grounded. Our framework advances the direction of agentic AI by enabling systems to adapt to user identity, intent, and context. By capturing surgeon-specific behavioral fingerprints and reasoning over structured gesture sequences, we take a step toward building surgical agents capable of individualized learning, proactive assistance, and trust-aware interaction in clinical environments.